# PITT-PACC-2408

### v1

@@ Appendix A Steps of the Kinematic Approach
where $g_{A}$ and $g_{V}$ are the vector and axial couplings of $f\bar{f}$ currents to the $Z$ boson, with $g_{V}=-1/4+\sin^{2}\theta_{W}$ and $g_{A}=-1/4$ for $e^{+}e^{-}$, $\mu^{+}\mu^{-}$, or $\tau^{+}\tau^{-}$.
@@
When $\beta=1$ this reduces to
$C_{ij}=\begin{pmatrix}1&0&0\\ 0&\frac{s_{\Theta}^{2}(g_{A}^{2}-g_{V}^{2})}{(1+c_{\Theta}^{2})(g_{A}^{2}+g_{V}^{2})}&0\\ 0&0&-\frac{s_{\Theta}^{2}(g_{A}^{2}-g_{V}^{2})}{(1+c_{\Theta}^{2})(g_{A}^{2}+g_{V}^{2})}\end{pmatrix}.$ (32)
@@
$F_{\mathcal{C}}(\Theta,\beta=1)$ $=\frac{1}{2}(C_{22}-C_{33}+C_{11}-1)=\frac{s_{\Theta}^{2}(g_{A}^{2}-g_{V}^{2})}{(1+c_{\Theta}^{2})(g_{A}^{2}+g_{V}^{2})},$ (34)
$F_{\mathcal{B}}(\Theta,\beta=1)$ $=\sqrt{2}(C_{11}+C_{22})=\frac{2\sqrt{2}\left(g_{A}^{2}+g_{V}^{2}c_{\Theta}^{2}\right)}{\left(1+c_{\Theta}^{2}\right)\left(g_{A}^{2}+g_{V}^{2}\right)}.$ (35)

### v2

@@ Appendix A Steps of the Kinematic Approach
where $\theta_{W}$ is the Weinberg angle, $g_{A}$ and $g_{V}$ are the vector and axial couplings of $f\bar{f}$ currents to the $Z$ boson, with $g_{V}=-1/4+\sin^{2}\theta_{W}$ and $g_{A}=-1/4$ for $e^{+}e^{-}$, $\mu^{+}\mu^{-}$, or $\tau^{+}\tau^{-}$.
@@
When $\beta=1$, both the cross section and the correlation matrix reduce to a simple form:
$\frac{1}{\sigma}\frac{d\sigma}{d\cos\Theta}=\frac{3}{8}(1+\cos^{2}\Theta),~{}~{}C_{ij}=\begin{pmatrix}1&0&0\\ 0&\frac{s_{\Theta}^{2}(g_{A}^{2}-g_{V}^{2})}{(1+c_{\Theta}^{2})(g_{A}^{2}+g_{V}^{2})}&0\\ 0&0&-\frac{s_{\Theta}^{2}(g_{A}^{2}-g_{V}^{2})}{(1+c_{\Theta}^{2})(g_{A}^{2}+g_{V}^{2})}\end{pmatrix}.$ (32)
@@
The full quantum tomography for $Z\to\mu^{+}\mu^{-}$ at the LHC is shown in Fig. 3.
Figure 3: Complete quantum tomography distribution for $Z\to\mu^{+}\mu^{-}$ at the LHC. All the other parameters that are not shown are zero.
@@
$\mathcal{C}(\Theta,\beta=1)$ $=\frac{1}{2}(C_{22}-C_{33}+C_{11}-1)=\frac{s_{\Theta}^{2}(g_{A}^{2}-g_{V}^{2})}{(1+c_{\Theta}^{2})(g_{A}^{2}+g_{V}^{2})},$ (34)
$\mathcal{B}(\Theta,\beta=1)$ $=\sqrt{2}(C_{11}+C_{22})=\frac{2\sqrt{2}\left(g_{A}^{2}+g_{V}^{2}c_{\Theta}^{2}\right)}{\left(1+c_{\Theta}^{2}\right)\left(g_{A}^{2}+g_{V}^{2}\right)}.$ (35)
and their variances $\mathrm{Var}(\mathcal{R})=\left\langle\left(\mathcal{R}-\overline{\mathcal{R}}\right)^{2}\right\rangle=\left\langle\mathcal{R}^{2}\right\rangle-\overline{\mathcal{R}}^{2}$ are calculated as follows,
$\overline{\mathcal{C}}$ $=\frac{1}{\sigma}\int_{-1}^{1}\frac{d\sigma}{d\cos\Theta}\mathcal{C}(\Theta,\beta=1)~{}d\cos\theta=\frac{g_{V}^{2}-g_{A}^{2}}{2(g_{V}^{2}+g_{A}^{2})}$ (36a)
$\braket{\mathcal{C}^{2}}$ $=\frac{1}{\sigma}\int_{-1}^{1}\frac{d\sigma}{d\cos\Theta}\left(\mathcal{C}(\Theta,\beta=1)\right)^{2}~{}d\cos\theta=\frac{3\pi-8}{4}\frac{(g_{V}^{2}-g_{A}^{2})^{2}}{(g_{V}^{2}+g_{A}^{2})^{2}}$ (36b)
$\Longrightarrow~{}\mathrm{Var}(\mathcal{C})$ $=\frac{3(\pi-3)}{4}\frac{(g_{V}^{2}-g_{A}^{2})^{2}}{(g_{V}^{2}+g_{A}^{2})^{2}}$ (36c)
and
$\overline{\mathcal{B}}$ $=\frac{1}{\sigma}\int_{-1}^{1}\frac{d\sigma}{d\cos\Theta}\mathcal{B}(\Theta,\beta=1)~{}d\cos\theta=\frac{g_{V}^{2}+3g_{A}^{2}}{\sqrt{2}(g_{V}^{2}+g_{A}^{2})}$ (37a)
$\braket{\mathcal{B}^{2}}$ $=\frac{1}{\sigma}\int_{-1}^{1}\frac{d\sigma}{d\cos\Theta}\left(\mathcal{B}(\Theta,\beta=1)\right)^{2}~{}d\cos\theta=\frac{3\pi g_{A}^{4}+6(4-\pi)g_{V}^{2}g_{A}^{2}+(3\pi-8)g_{V}^{4}}{2(g_{V}^{2}+g_{A}^{2})^{2}}$ (37b)
$\Longrightarrow~{}\mathrm{Var}(\mathcal{B})$ $=\frac{3(\pi-3)}{2}\frac{(g_{V}^{2}-g_{A}^{2})^{2}}{(g_{V}^{2}+g_{A}^{2})^{2}}=2\mathrm{Var}(\mathcal{C}).$ (37c)